# Market Reaction to News Flows in Supply Chain Networks

### v1

@@ Abstract
This study examines whether positive news about firms increases their stock prices and, moreover, whether it increases stock prices of the firms’ suppliers and customers, using a large sample of publicly listed firms across the world and another of Japanese listed firms. The level of positiveness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys. We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through informal channels. Positive news also raises stock prices of the firms’ suppliers and customers before its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and customers than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for informal information diffusion. However, the post-news effect on suppliers and customers in Japan is smaller than the pre-news effect, a result opposite to those from firms across the world. This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
Introduction
Firms interact with and influence each other through various types of networks. Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 2, 3] and knowledge networks through research collaboration.[4, 5, 6, 7, 8] Another type of network recently paid attention in the academic literature and business practices is supply chains formed through transactions of materials, parts, and components.[9, 10] One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other. [11]
@@ Introduction
Therefore, positive or negative shocks propagate through supply chains, affecting performance of firms linked through supply chains. Because the corporate value of publicly listed firms is evaluated in the stock market, [24, 25, 26] a shock to a firm is expected to affect not only its own stock price but also the stock price of its suppliers and customer firms. For example, when Boeing 737 MAX crashed in 2019 and 2020, the stock price of Boeing, its supplier, such as GE and Allegheny Technologies, and its customers, such as American Airlines and Southwest Airlines, dropped sharply.[27, 28, 29, 30, 31]. Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 negatively affected the stock price of Daihatsu, Toyota, and their suppliers and customers, such as Aisin.[32, 33]
@@
These examples show that the shock can propagate through supply chains and affect stock prices of firms linked through supply chains. Although previous studies have examined the direct impact of shocks to firms on their own stock prices[34, 35, 36], no study has investigated the diffusion of shocks to other firms’ stock prices through supply chains to the best of our knowledge. To fill the gap, this study examines this possible diffusion, using two large firm-level samples, one for firms across the world and the other for Japanese firms. Specifically, we test whether the level of negative sentiment about a firm in a news article has a positive impact on the change rate of the stock price of the firm and its suppliers and customers before and after the disclosure of the news.
@@
We contribute to the literature by utilizing large-scale data in many respects. First, our firm-level data covers all publicly listed firms in the world and in Japan and identify detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys. Second, our data on news articles are taken from the NewsArchive of Thomson Reuters and cover more than 20 million articles. Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive news. Another notable contribution of this study is that we utilize sentiment analysis to determine the positiveness of each news article. Prior research on sentiment analysis has predominantly focused on neural networks [37, 38], and recently, Bidirectional Encoder Representations from Transformers (BERT)[39] has garnered the most acclaim as the state-of-the-art approach [35]. We take advantage of FinBERT,[40] an extended model of BERT particularly for financial information.
Data
@@
This study utilizes four major data sources. First, our information on news relies on news articles written in English included in the NewsArchive provided by Thomson Reuters for the period 2003-2016. The information in each news article includes the date, headline, main text, International Securities Identification Numbers (ISIN) of firms mentioned in the article, and language. As we will explain later in detail, our sentiment analysis to determine whether each news article is positive or negative about firms mentioned there utilizes the main text. The average number of words in an article is 243, whereas its median is 136 and maximum is 3911. Although the total number of news articles in the NewsArchve for the period examined is 20,803,561, we focus on 3,447,425 news articles that mention any firm publicly listed in 105 stock markets in the world. 90% of these news articles mention 3 or fewer firms, whereas the maximum number of mentioned firms in an article is 33 (Figure 1). The number of news articles that mention a particular firm during the period 2003-2016 is quite skewed. **% of firms were never mentioned, and **% were mentioned only once, while 0.1% were mentioned more than 10,000 times (Figure 2).
@@
Second, we utilize daily stock prices of currently listed firms in stock markets in the world taken from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, using its Application Programming Interface (API). We drop firms listed in the past but not currently from our sample, because their stock prices are not available using the API. We particularly define the daily stock price of each firm as its closing price. In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average of stock prices in the market to control for the market-specific time trend. The Refinitiv index is available for 51 major markets that cover 92% of listed firms worldwide.
@@
Third, to identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003-2016 obtained as of January 18, 2023. The FactSet data include major clients of each firm reported in its financial statement. The US, Japanese, and many other governments require every listed firm to disclose its major clients with which the sales are more than 10% of its total sales, at least in the period examined in this study [41]. Therefore, based on published financial statements and, in addition, supplemental information from news articles and web sites, FactSet identifies supply chains of listed and non-listed firms with their major partners. Table 1 shows the number of all firms (for reference) and listed firms (our sample) in the FactSet data and their supply-chain links by year. The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed more recently and thus ignores firms in the distant past.
@@
| firms across the world (listed) | |||||||
@@
| firms across the world (listed) | |||||||
@@
Finally, to identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008-2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan. Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms of each firm. Apparently, many firms are linked with more than 23 suppliers and clients, and thus their links cannot be fully identified by their own responses. However, these links can be mostly identified by responses by their suppliers and clients. Accordingly, the maximum number of suppliers and clients is *** and ***, respectively, in 2016, implying that our data cover supply chains of Japanese firms to a great extent. The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
| 2008 | 2009 | 2010 | 2011 | 2012 | 2013 | 2014 | 2015 | 2016 | |
| Japan (total) | |||||||||
| Number of firms | 1,018,570 | 1,061,051 | 1,107,760 | 1,135,200 | 1,156,205 | 1,182,729 | 1,194,034 | 1,200,162 | 1,247,939 |
| Number of links | 4,443,677 | 4,644,977 | 4,810,967 | 4,925,307 | 4,995,612 | 5,090,119 | 5,157,181 | 5,209,604 | 5,491,417 |
| Japan (listed firms) | |||||||||
| Number of firms | 3,554 | 3,456 | 3,403 | 3,308 | 3,245 | 3,215 | 3,204 | 3,215 | 3,205 |
| Number of links | 21,720 | 22,128 | 22,630 | 22,044 | 21,833 | 21,696 | 21,439 | 20,968 | 20,445 |
Methods
@@
Our sentiment analysis to determine the degree of positiveness and negativeness of the information about the listed firms provided in each news article utilizes a natural language processing (NLP) model called FinBERT[40]. FinBERT is an extended version of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al. (2018)[39]. Specifically, FinBERT uses Reuters’ TRC2-financial that consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank that consists of 4,845 English sentences taken from financial news in the LexisNexis database. From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the probabilities of the three about the firms mentioned in the article. Before conducting the sentiment analysis using FinBERT, we clean news articles by deleting URLs, line breaks, ISINs, and fixed phrases at the end.
The distribution of the probabilities of positive, neutral, and negative sentiments generated by FinBERT from news articles about firms across the world and Japanese firms are provided in Panels (A) and (B) of Figure 3, respectively. Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.
From the probabilities of the three types of sentiment, we construct an index of positive sentiment by the ratio of the probability of positive sentiment to the sum of the probabilities of positive and negative sentiment, ignoring the probability of neutral sentiment. For example, when the probabilities of positive, neutral, and negative sentiment are 0.3, 0.5, and 0.2, respectively, our positive sentiment index is . We ignore the probability of neutral sentiment to avoid the ambiguity of sentiment in news articles. Figure 4 demonstrates the distribution of this positive sentiment index for firms across the world and Japanese firms. Because we avoid the ambiguity of the sentiment, the distribution of the index is skewed to both lower (0) and upper (1) limits although we observe a hike around the value of 0.25. In other words, our positive sentiment index distinguishes between positive and negative sentiment more clearly than the probabilities of the three directly generated by FinBERT and highlights positive sentiment in news articles.
Regression analysis
@@
To estimate the effect of positive sentiment in news articles about particular firms on their own stock prices, we apply the positive sentiment index constructed above to regression analysis. Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window are affected by the level of positiveness of the article after the article is disclosed, compared with its change rate before the disclosure. Suppose that firm listed in stock market is mentioned by news article on day . Setting the time window at , such as 2, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article disclosed on day in time window , or and , respectively, by
@@
where is the average stock price between day and . For example, is the average stock price of firm for days between day and . Therefore, the left-hand side of equations (1) and (2) indicates the daily percentage change of the average stock price in time window in the pre- and post-news period, respectively. We take the average of stock prices over the time window to reduce their fluctuations. We exclude Saturdays and Sundays from our daily observations because stock prices are not available on weekends. Accordingly, for example, when news article about firm is disclosed on Thursday, the 10th of June and the time window is set at 3 days, the pre-news change rate is defined as the change rate from the average stock price on June 3rd (Thursday), 4th (Friday), and 7th (Monday), skipping 5th (Saturday) and 6th (Sunday), to the average on 8th (Tuesday)-10th (Thursday), as shown in Figure 5. Similarly, the post-news change is defined as the change from the average on June 7th-9th to that on June 10th, 11th (Friday), and 14th (Monday).
@@
Thus, our sample is restricted to firm-day observations before or after news articles about the firms are disclosed. In other words, the number of observations in our sample is the cumulative number of firms mentioned in news articles during the period examined multiplied by 2. Using the pre- and post-news observations, we estimate the following estimation equation:
@@
where and are indicator variables that take a value of one if is the percentage change of the stock price of firm before and after news article is disclosed on day , respectively, or . represents the level of positive sentiment in news article generated by FinBERT (Data section). is the percentage change of the Refinitiv index for each stock market that controls for the overall time trend of stock prices in the market, and is sector dummies. Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industry Classification at the two-digit level for the sample of Japanese firms.
@@
Accordingly, indicates the average effect of the positiveness of a news article on the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors. By contrast, indicates any prior “effect” of the news article on the change rate of the stock price, including the possible effect of the information in the news article that diffuses in the market not through the news but through other informal channels before the disclosure of the news. In addition, may include the endogeneity bias reflecting the possibility that the growth rate of the stock price of a firm that is positively mentioned by any news article is intrinsically higher than that of a firm negatively mentioned. Therefore, the net effect of the disclosure of the information by the news article on the change rate of the stock price in time window should be given by [42].
@@
We particularly estimate equation (3) for to examine both the short- and long-run effect of positive news articles on stock prices. We hypothesize that a positive news article would accelerate the change rate of the stock price of the firm mentioned in the article in the short-run, while the effect would become smaller and converge to 0 in the longer run. Note that equation (3) for cannot be estimated because the change rates of the stock price in the pre- and post-news periods are the same, according to equations (1) and (2).
@@
We further examine the effect of positive sentiment about firms in news articles on their suppliers and customers, using similar frameworks. Specifically, when we focus on the upstream effect of news articles about firms on their suppliers, in equation (3) is replaced with the probability of positive sentiment in the news article that mentions not firm but any of its customer firms. Alternatively, when we examine the downstream effect of customers’ news on their customers, is replace with the positive index of any of suppliers of firm . When a firm mentioned by a news article has more than one supplier or customer, as shown in Tables 1 and 2, we include all the suppliers and customers in our sample. In this analysis, we compare stock prices of suppliers or customers of firms mentioned positively in news articles with those of supplier or customers of firms mentioned negatively, as in our analysis on the direct effect on own stock prices.
We hypothesize that the effect of news articles on stock prices diffuses through supply chains. Precisely, a positive news article about a customer firm would accelerate the growth in the stock price of its suppliers, in addition to its own stock price, because the market predicts an increase in demand for products of the suppliers associated with higher performance in the customer firm. Also, positive news about a supplier would promote growth of its customers because of possible productivity growth of the supplier. As in the case of the direct effect, the indirect effect on suppliers and customers would be smaller in the long run than in the short run.
Results and Discussion

### v3

@@ Abstract
This study examines how positive and negative news about firms affects their stock prices and, moreover, how it affects stock prices of the firms’ suppliers and clients, using a large sample of publicly listed firms around the world and another of Japanese listed firms. The level of positiveness and negativeness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys. We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through unofficial information channels. Positive news also raises stock prices of the firms’ suppliers and clients before and after its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and clients than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for information diffusion through private channels. However, the post-news effect on suppliers and clients in Japan is smaller than the pre-news effect, which is the opposite result to non-domestic firms from around the world.
keywords:
Supply Chain, Natural Language Processing, Stock Market, Large -scale Data1 Introduction
Firms interact with and influence each other through various types of networks. Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 2, 3] and knowledge networks through research collaboration.[4, 5, 6, 7, 8] Another type of network recently paid attention in the academic literature and business and policy fields is supply chains formed through transactions of materials, parts, and components.[9, 10] One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other. [11]
@@
Therefore, positive or negative shocks propagate through supply chains, affecting performance of firms linked through supply chains. Because the corporate value of publicly listed firms is evaluated in the stock market,[24, 25, 26] a shock to a firm is expected to affect not only its own stock price but also the stock prices of its suppliers and client firms. For example, when Boeing 737 MAX crashed in 2019 and 2020, the stock price of Boeing, its supplier, such as GE and Allegheny Technologies, and its clients, such as American Airlines and Southwest Airlines, dropped sharply.[27, 28, 29, 30, 31]. Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 negatively affected the stock price of Daihatsu, Toyota, and their suppliers and clients, such as Aisin.[32, 33]
These examples show that the shock can propagate through supply chains and affect stock prices of firms linked through supply chains. Although relatively a lot of previous studies have examined the direct impact of shocks to firms on their own stock prices[34, 35, 36], there are not so many studies that investigated the diffusion of shocks to other firms’ stock prices through supply chains. Kim and Wagner [37] analyzed 315 cases reported by the Wall Street Journal regarding supply chain risks, which also included mentions of 69 suppliers and 246 clients. Pandit et al. utilized a larger dataset of quarterly earnings announcements, which differ from the news reports considered in our study, and analyzed their propagation across 1,518 suppliers and 4,010 clients in COMPUSTAT[38]. In addition, similar studies focusing on China are discussed [39, 40]. However, to the best of our knowledge, no study has examined large-scale data spanning more than a decade and covering tens of thousands of firms.
@@
To fill this gap, this study examines this possible diffusion, using two large firm-level samples, one for firms across the world and the other for Japanese firms. Specifically, we test whether the level of positive and negative sentiment about a firm in a news article affects the change rate of the stock price of the firm and its suppliers and clients before and after the disclosure of the news.
@@
We contribute to the literature by utilizing large-scale data in many respects. First, our firm-level data covers most publicly listed firms in the world and in Japan and identify detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys. Second, our data on news articles are taken from the NewsArchive of Thomson Reuters and cover more than 20 million articles. Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive news.
Although we have mainly focused on how the shock affects stock prices through supply chains, it is evident that this study also relates to business applications and researches that predict changes of stock prices based on the news. In this field, neural networks are predominantly used. For example, hedge funds, such as Bridgewater Associates utilizes the neural networks for predictions[41]. In academia, the transformer[42] is widely used[43, 44]. In addition to the transformer, there is a long history of the usage of neural networks to predict stock prices, such as long short-term memory[45, 46, 47, 48, 49, 50, 51, 52], convolutional neural networks[53], or other deep learning methods[54, 55, 56, 57]. We use sentiment analysis to determine the positiveness and negativeness of each news article[58, 59] and for this purpose, specifically, Bidirectional Encoder Representations from Transformers (BERT)[60] has garnered the most acclaim[35]. We take advantage of FinBERT[61], a fine-tuned model of BERT particularly for financial information.
2 Data
@@
This study utilizes four major data sources. First, our information on news relies on news articles written in English included in the NewsArchive provided by Thomson Reuters for the period 2003-2016. The information in each news article includes the time, headline, main text, Refinitiv Identification Code (a ticker-like code) of firms mentioned in the article, language, etc. As we will explain in detail later, our sentiment analysis, which determines the degree of positivity or negativity of each news article about the firms mentioned, utilizes the main text. The average number of words in an article is 243, whereas its median is 136 and maximum is 3,911. Although the total number of news articles in the NewsArchive for the period examined is 20,803,561, we focus on 3,447,425 that mention any firm publicly listed in 105 stock markets in the world. Among these news articles, 90% mention three or fewer firms, while the maximum number of firms mentioned in a single article is 33. (Figure 1). The number of news articles that mention a particular firm during the period 2003-2016 is quite skewed. 55.6% of firms were never mentioned, while 0.1% were mentioned more than 10,000 times (Figure 2).
@@
Second, we utilize daily stock prices of currently listed firms in stock markets in the world taken from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, using its Application Programming Interface (API). We drop firms listed in the past but not currently from our sample, because their stock prices are not available using the API. We particularly define the daily stock price of each firm as its closing price. In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average of stock prices in the market to control for the market-specific time trend. The Refinitiv index is available for 105 major markets that cover 92% of listed firms worldwide.
@@
Third, to identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003–2016 obtained as of January 18, 2023. The FactSet data include major clients of each firm reported in its financial statement. The US, Japanese, and many other governments require every listed firm to disclose its major clients with which the sales are more than 10% of its total sales, at least in the period examined in this study [62]. Therefore, based on published financial statements and, in addition, supplemental information from news articles and web sites, FactSet identifies supply chains of listed and non-listed firms with their major partners. Table 1 shows the number of all firms (for reference) and listed firms (our sample) in the FactSet data and their supply-chain links by year. The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed recently and thus ignores firms in the distant past. In a later section, the recorded supply chain links at the time are considered to estimate the impact of the news on supply chains.
@@
| Maximum indegree | 283 | 272 | 272 | 421 | 343 | 291 | 195 |
| Maximum outdegree | 255 | 266 | 249 | 203 | 159 | 180 | 156 |
| Listed | |||||||
@@
| Maximum indegree | 463 | 600 | 577 | 496 | 423 | 586 | 562 |
| Maximum outdegree | 162 | 201 | 206 | 198 | 256 | 294 | 378 |
| Listed | |||||||
@@
| Maximum indegree | 27 | 70 | 115 | 127 | 161 | 225 | 277 |
| Maximum outdegree | 40 | 54 | 79 | 79 | 75 | 109 | 140 |
Finally, to identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008–2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan. Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms of each firm. Apparently, many firms are linked with more than 23 suppliers and clients, and thus their links cannot be fully identified by their own responses. However, these links can be mostly identified by responses by their suppliers and clients. Since those firms are relatively large, their partners report them as important suppliers or clients. The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
| 2008 | 2009 | 2010 | 2011 | 2012 | |
| All firms | |||||
| Number of firms | 1,018,570 | 1,061,051 | 1,107,760 | 1,135,200 | 1,156,205 |
| Number of links | 4,443,677 | 4,644,977 | 4,810,967 | 4,925,307 | 4,995,612 |
| Maximum indegree | 6,751 | 6,976 | 7,069 | 7,332 | 8,398 |
| Maximum outdegree | 11,140 | 11,150 | 11,155 | 11,201 | 11,287 |
| Listed | |||||
| Number of firms | 3,554 | 3,456 | 3,403 | 3,308 | 3,245 |
| Number of links | 21,720 | 22,128 | 22,630 | 22,044 | 21,833 |
| Maximum indegree | 201 | 189 | 180 | 162 | 151 |
| Maximum outdegree | 305 | 277 | 259 | 243 | 225 |
| 2013 | 2014 | 2015 | 2016 | |
| All firms | ||||
| Number of firms | 1,182,729 | 1,194,034 | 1,200,162 | 1,247,939 |
| Number of links | 5,090,119 | 5,157,181 | 5,209,604 | 5,491,417 |
| Maximum indegree | 8,787 | 9,520 | 9,948 | 12,016 |
| Maximum outdegree | 11,490 | 11,586 | 11,588 | 12,729 |
| Listed | ||||
| Number of firms | 3,215 | 3,204 | 3,215 | 3,205 |
| Number of links | 21,696 | 21,439 | 20,968 | 20,445 |
| Maximum indegree | 148 | 145 | 143 | 141 |
| Maximum outdegree | 216 | 201 | 195 | 190 |
3 Methods
@@
Our sentiment analysis to determine the degree of positiveness and negativeness of the information about the listed firms provided in each news article utilizes a natural language processing (NLP) model called FinBERT[61]. FinBERT is a variant of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al. (2018)[60]. Specifically, FinBERT uses Reuters’ TRC2-financial that consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank that consists of 4,845 English sentences taken from financial news in the LexisNexis database. From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the weights of the three about the firms mentioned in the article. Examples are shown in Table 3. Before conducting the sentiment analysis using FinBERT, we clean news articles by deleting URLs, line breaks, ISINs, and fixed phrases at the end. The distribution of the probabilities of positive, neutral, and negative sentiments generated by FinBERT from news articles about firms across the world and Japanese firms are provided in Panels (A) and (B) of Figure 3, respectively. Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.
| Sentiment | |||
| News article | Positive | Neutral | Negative |
| (Sep/11/2008) Advanced Medical Solutions Group Plc on Thursday said its silver anti-microbial wound gel had been approved by the U.S. Food & Drug Administration (FDA), sending its shares up 5 percent. … [Source: [63]] | 0.93 | 0.01 | 0.06 |
| (Sep/11/2008) U.S. Interior Department employees who oversaw oil drilling on federal lands had sex and used illegal drugs with workers at energy companies … [Source: [64]] | 0.02 | 0.21 | 0.76 |
| (Sep/11/2008) Amazon.com, the largest global online retailer, plans to start selling U.S.-produced wine on its website within the United States by late September or October … [Source: [65]] | 0.14 | 0.84 | 0.01 |
3.1 Regression analysis
@@
To estimate the effect of positive and negative sentiment in news articles about particular firms on their own stock prices, we apply the probabilities of positive and negative sentiment constructed above to regression analysis. Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window are affected by the level of positiveness or negativeness of the article after the article is disclosed, compared with its change rate before the disclosure. Suppose that firm listed in stock market is mentioned by news article on day . Setting the time window at , such as 1, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article disclosed on day in time window , or and , respectively, by
@@
where is the average stock price of firm between day and . For example, is the average stock price of firm for days between day and . Therefore, the left-hand side of equations (1) and (2) indicates the daily percentage change of the average stock price in time window in the pre- and post-news period, respectively. We take the average of stock prices over the time window to reduce their fluctuations. We skip days when stock prices are unavailable, such as Saturdays, Sundays, and holidays. Accordingly, for example, as shown in Figure 4, when news article about firm is disclosed on Friday, the 11th of June and the time window is set at 3 days, the pre-news change rate is defined as the change rate from the average stock price on June 3rd (Thursday), 4th (Friday), and 7th (Monday), skipping 5th (Saturday) and 6th (Sunday), to the average from the 8th (Tuesday) to 10th (Thursday). Similarly, the post-news change is defined as the change from the average of June 8–10 to that of June 11th, 14th (Monday), and 15th (Tuesday).
@@
Thus, our sample is restricted to firm-day observations that occur before or after news articles about the firms are disclosed. In other words, the number of observations in our sample is the cumulative number of firms mentioned in news articles during the period examined multiplied by 2. Using the pre- and post-news observations, we estimate the following estimation equation:
@@
where T is either pre or post, and and are indicator variables that take a value of one if represents the percentage change in the stock price of firm before or after news article is disclosed on day , respectively. represents the level of either positive or negative sentiment in news article generated by FinBERT (Data section). is the percentage change of the Refinitiv index for each stock market that controls for the overall time trend of stock prices in the market, and is sector dummies. Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industrial Classification at the two-digit level for the sample of Japanese firms.
@@
Accordingly, indicates the average effect of the positiveness or negativeness of a news article on the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors. By contrast, indicates any prior “effect” of the news article on the change rate of the stock price, including the possible effect of the information in the news article that diffuses in the market not through the news but through unofficial information channels before the disclosure of the news. Such unofficial information channels include suggestions by financial companies to their clients, posts in Social Networking Services (SNS) by influencers, and word of mouth among investors. In addition, may include the endogeneity bias reflecting the possibility that the growth rate of the stock price of a firm that is positively mentioned by any news article is intrinsically higher than that of a firm negatively mentioned. Therefore, the net effect of the disclosure of the information by the news article on the change rate of the stock price in time window should be given by [66].
@@
We particularly estimate equation (3) for 1, 2, 3, 4, 5, 30, 180, and 365 to examine both the short- and long-run effect of positive news articles on stock prices. We hypothesize that a positive (negative) news article would accelerate (deteriorate) the change rate of the stock price of the firm mentioned in the article in the short-run, while the effect would become smaller and converge to 0 in the longer run.
@@
We further examine the effect of positive or negative sentiment about firms in news articles on their suppliers and clients, using similar frameworks. Specifically, when we focus on the upstream effect of news articles about firms on their suppliers, in equation (3) is replaced with the probability of positive or negative sentiment in the news article that mentions not firm but any of ’s client firms. Alternatively, when we examine the downstream effect of clients’ news on their clients, is replaced with the positive or negative index of any of the suppliers of firm . When a firm mentioned by a news article has more than one supplier or client, as shown in Tables 1 and 2, we include all the suppliers and clients in our sample. In this analysis, we compare stock prices of suppliers or clients of firms mentioned positively in news articles with those of suppliers or clients of firms mentioned negatively, as in our analysis on the direct effect on own stock prices.
We hypothesize that the effect of news articles on stock prices diffuses through supply chains. Precisely, a positive news article about a client firm would accelerate the growth in the stock price of its suppliers, in addition to its own stock price, because the market predicts an increase in demand for products of the suppliers associated with higher performance in the client firm. Also, positive news about a supplier would promote growth of its clients because of possible productivity growth of the supplier. As in the case of the direct effect, the indirect effect on suppliers and clients would be smaller in the long run than in the short run.
4 Results and Discussion